# Analyzing News Engagement on Facebook: Tracking Ideological Segregation and News Quality in the Facebook URL Dataset

### v2

@@ Abstract
The Facebook Privacy-Protected Full URLs Dataset was released to enable independent, academic research on the impact of Facebook’s platform on society while ensuring user privacy. The dataset has been used in several studies to analyze the relationship between social media engagement and societal issues such as misinformation, polarization, and the quality of consumed news. In this paper, we conduct a comprehensive analysis of the engagement with popular news domains, covering four years from January 2017 to December 2020, with a focus on user engagement metrics related to news URLs in the U.S. By incorporating the ideological alignment and quality of news sources, along with users’ political preferences, we construct weighted averages of ideology and quality of news consumption for liberal, conservative, and moderate audiences. This allows us to track the evolution of (i) the ideological gap in news consumption between liberals and conservatives and (ii) the average quality of each group’s news consumption. We identify two major shifts in trends, each tied to engagement changes. In both, the ideological gap widens and news quality declines. However, engagement rises in the first shift but falls in the second. Finally, we contextualize these trends by linking them to two major Facebook News Feed updates. Our findings provide empirical evidence to better understand user behavior, polarization, and misinformation during the period covered by the dataset.
@@ 1 Introduction
The interplay between user ideology, news source quality, and online engagement has been a key focus of research in recent years. Studies consistently show that individuals tend to consume ideologically aligned content, with less moderate views often linked to lower-quality news sources Bakshy et al. (2015); Robertson et al. (2023), and more extreme narratives being associated with higher engagement Galeazzi et al. (2024). Extreme conservative audiences, in particular, have been found to exhibit higher levels of ideological segregation González-Bailón et al. (2023); Heseltine (2023). These patterns raise concerns about the relationship between online platforms’ design and the segregation of information consumption, as platform-specific mechanisms can influence the extent of ideological separation. For instance, ideological segregation appears stronger on Facebook than on Reddit, where interactions remain more heterogeneous even among extreme-leaning users Cinelli et al. (2021). Moreover, selective exposure to partisan news sources tends to increase with user activity Cinelli et al. (2020), contributing to the formation of highly clustered news consumption patterns Schmidt et al. (2017). Higher engagement within these echo chambers has also been associated with a shift toward more negative emotional expression, which may further shape user interactions and content dynamics Del Vicario et al. (2016).
@@
Building on the literature on algorithmic amplification, recommender systems have been shown to prioritize engagement, often at the expense of exposing users to more polarizing or toxic content, thereby amplifying ideological divides Chavalarias et al. (2024). Regarding the assessment of partisan audience bias in social media, Robertson et al. (2018, 2023) have conducted audits and investigations on user exposure and interactions with content on Google Search, while Bakshy et al. (2015) used US Facebook data on interactions with shared news and friend networks. Contrarily to González-Bailón et al. (2023), both concluded that individual choices played a stronger role in increasing like-minded content, as well as limiting cross-cutting exposure than algorithmic ranking. These contrasting findings suggest the need for a more comprehensive analysis of user behaviors and algorithmic influences Budak et al. (2024).
@@
The Facebook Privacy Protected Full URLs Dataset Messing et al. (2020) is a large-scale collection of public posts containing URLs and user interaction data on the Facebook platform starting January 2017. The dataset contains interaction counts of engagement at the URL-action level, with some stratification on the information of users who issue those actions (i.e. political page affinity, age, and gender). It does not contain, however, information on the volume of unique users in each of these categories or in the dataset in general. This is a dataset unique in its size and richness but has its peculiarities regarding privacy-protecting methods and limitations, which we address in Section 2.2 and Supplementary Information 1A-D.
@@
Despite these limitations, researchers have used the Facebook Privacy-Protected Full URLs Dataset for various applications. Our work, akin to Guess et al. (2021) and Bailey et al. (2021), provides a descriptive static analysis of exposure and sharing patterns of news URLs in the US, confirming findings like the conservative skew in low-quality news consumption. Building on this foundation, our work introduces a temporal dimension to investigate ideological segregation and information quality in news consumption. Similar to Buntain et al. (2023), who developed a metric for robust domain ideology scores based on exposure and engagement, we focus on user behavior and content interactions but expand the scope by examining longitudinal trends and dynamic patterns over time. Other works have used the dataset for information extraction Allen et al. (2024); Barnabò et al. (2022, 2023); Heuer and Glassman (2022).
@@
Related research on engagement with misinformation, such as the studies by Thero and Vincent (2022) and Vincent et al. (2022), highlights the limited long-term effectiveness of labeling and penalizing pages tagged as “repeat offenders”. These findings stress the complexity of combating misinformation, an issue we approach by analyzing how engagement with low-quality news changes over time and how ideological leanings influence these patterns. Bandy and Diakopoulos (2023) show that Facebook amplified low-quality publishers during the 2020 US election, with algorithmic changes impacting publishers broadly, regardless of content quality. We extend this line of inquiry by contextualizing engagement trends within potential algorithmic adjustments Mosseri (2018); Hagey and Horwitz (2021); Narayanan (2023) to examine their implications for ideological segregation and polarization.
@@
Our contributions focus on the evolution of engagement patterns with both news-related URLs and other types of content over a four-year period. Specifically, for popular news, we characterize engagement trends (e.g., clicks, likes, shares, and comments) in relation to user ideology, domain ideology, and domain quality. By developing a novel metric to assess partisan bias in user consumption patterns, we track the evolution of the ideological gap between news consumed by conservatives and liberals as a proxy for polarization in news diets. Additionally, we perform a detailed study of changes in the quality of news consumption and its relation to user and news outlet ideology.
@@
Our findings provide a detailed view of the dynamic relationship between user behavior, news quality, and ideological polarization in news diets, while situating these changes within the broader context of platform-level adjustments and potential algorithmic modifications.
@@ 2 Data Collection, Processing, and Descriptive Analysis
For the US, these counts are also stratified by the estimated ideological leaning of the users, called Political Page Affinity or PPA, which classifies the audience into five different buckets, scaled from $-2$ (most liberal) to $+2$ (most conservative), with an additional bin (not defined). This metric, based on the model described in Barberá (2015), is computed internally at Facebook. For that reason, we focus on URLs shared primarily in the US and extract engagement counts from users located in the US only.
@@ 2.1 Domain and Data Selection Criteria
For this study, we separate between: (i) popular news domains, relative to news content, which have both quality ratings and ideological scores, and (ii) other (potentially non-news) domains, which cover the rest of the domains included in the Facebook URL dataset.
The selection of news domains is done three-way. We first restrict our list to the $2,629$ top 1% most-viewed domains on Facebook between 2017 and 2021, which we take from Buntain et al. (2023). We then restrict the list of news-related domains to only those domains for which we have the following information:
@@
    Quality scores: we use the open-source rating set from Lin et al. (2023), which is highly correlated with other widely used rating systems, such as NewsGuard [^1] https://www.newsguardtech.com/. Both ratings use a scale from 00 to $1$. Among the $2,629$ domains, $1,586$ ($\approx 60\%$) are found in the rating set of Lin et al. (2023).
@@
    Ideology scores: we use the domain-level ideology scores from Robertson et al. (2018), which include $1,664$ of the $2,629$ domains ($\approx 63\%$). Alternative ideology scores, such as those from Buntain et al. (2023), have been derived by aggregating user ideology estimates for each URL via PPA, and therefore they are unsuitable for our analyses on the same dataset. Nonetheless, Buntain et al. (2023) show that their scores are highly correlated with those we use, as well as with scores from previous studies Bakshy et al. (2015); Eady et al. (2020); Budak et al. (2016). For robustness, we repeated our analyses using the Buntain et al. (2023) scores and found no significant differences (data not shown).
We consider as news-related domains (or simply news) the intersection of the three lists (top 1% most-viewed domains, domains with quality scores, and domains with ideology scores), resulting in $1,231$ domains, of which approximately $58\%$ are high-quality (i.e. quality score larger than $0.6$, as suggested by NewsGuard’s guidelines Lin et al. (2023)). Conversely, we consider the domains that are not in intersection, as non-news related (or simply other). See Table 1 for the number of domains and related URLs for each of the two categories.
@@
As Figure 1 shows, there exists a weak negative correlation between the domain ideology and domain quality scores. We can also observe two clusters of high and medium-high quality located near the center ideology and a low-quality cluster at the conservative extreme.
@@
Figure 1: Selected news domains for this study distributed by Ideology and Quality scores. Positive values of ideology denote a leaning toward conservative ideology, while negative values correspond to liberal leanings. Quality scores are higher when domains have higher quality. The two dimensions correlate negatively ($r=-0.35,p=1.8\cdot 10^{-37}$).
@@ 2.2 URL Retrieval and Preprocessing
We retrieve URLs through successive monthly queries to the platform, covering the period from January 2017 to December 2020. The dataset is protected by differential privacy Messing et al. (2020); Evans and King (2023), which prevents tracing actions to specific users or URLs. However, while differential privacy safeguards data, it can introduce biases into inference methods if not properly accounted for Evans and King (2023). Additional challenges of this dataset include selection bias caused by engagement thresholds for URL inclusion Allen et al. (2021) and past data integrity issues Allen et al. (2022). An example of a precise query with parameters is detailed in Supplementary Fig. S1, along with a discussion of these limitations and our strategies for addressing them.
@@
To justify our choice of a cutoff period of three months, we calculate the typical lifespan of a URL. Figure 2B shows how engagement decays (on average) over time for the year 2019 (all years behave similarly) for three different types of engagement. We compute the total engagement for a URL within that year and normalize the proportion of engagement for each month, starting from the month in which it was first posted and continuing with subsequent months. The curve illustrates the expected lifespan of a URL’s engagement once it appears in the dataset. We observe that most of the engagement ($\approx 70\%$) occurs during the first two months. We repeated all our analyses using different values for the cutoff period (3-month, 6-month, and 12-month cutoffs) and found no significant differences in the results (data not shown). This indicates that extending the cutoff period essentially increases the size of the monthly URL set (see 2A), but the overall impact on engagement is negligible. Consequently, we present our results considering engagement counts using a 3-month cutoff.
@@ 3 Aggregate Analysis of Engagement
First, we analyze the distribution of engagement across user PPA categories. Since the dataset is designed to include noisy aggregated interaction counts per PPA (without information about the number of unique users per interaction), we experimented with different baselines and normalization methods. Details on the methods used to compute confidence intervals for engagement counts can be found in the Supplementary Information 2A.
@@
Normalizing by views captures the engagement given the exposure to the content and normalizing by the maximum value allows to compare across different types of engagement, despite their varying volumes. Alternative normalization methods and resulting outcomes are discussed in the Supplementary Information 3.
@@
Figure 3: Aggregate Analysis of Engagement (for popular news domains). In all figures, negative values correspond to more liberal-leaning, and positive values correspond to more conservative-leaning. (A) Distribution of different engagement types across user political page affinity (PPA), normalized relative to view counts (top left). For engagement types other than views, values are further normalized relative to their maximum across the five PPA categories, see Eq. (1). Engagement is unevenly distributed, with liberals having significantly more views than conservatives. The U-shaped pattern of engagement (greater engagement by extreme users) becomes more pronounced from left to right and from top to bottom. Users with undefined PPA exhibit much higher passive engagement (views and clicks) but significantly fewer engagement counts per view and are therefore not represented. (B) Conditional probability of clicking on a URL with a specific ideology given the user PPA. (bin width $=0.094$, columns sum one). Liberal and center users follow single-mode distributions of engagement, with liberals leaning towards negative (liberal) domain scores and center users favoring neutral domains. Conservative users display a bimodal distribution, engaging both with neutral and far-right domains. Undefined users exhibit a profile similar to center users. (C) Conditional probability of clicking on a URL with a specific quality given the user PPA. (bin width $=0.05$, columns sum one). Liberal users predominantly engage with high-quality domains. Center and conservative users progressively interact with lower-quality domains, with extreme conservatives showing the highest engagement with low-quality domains. This trend is supported by a negative correlation between domain quality and user ideology. See main text for details.
@@
Figure 3A shows the results for popular news domains. Engagement is unevenly distributed, with liberals having significantly more views than conservatives. This suggests that liberals tend to engage more frequently in passive consumption of news-related URL-containing posts, which we find to be opposite to other content (see Supplementary Fig. S4). For clicks, the other type of passive engagement, the pattern shifts slightly to a U-shape, with center users having the lowest number of clicks. As for the active engagement counts, there is a clear U-shaped pattern for hahas, likes, comments, and angers (increasingly pronounced in that same order). Surprisingly, the distribution for shares is the less U-shaped of all, with other reactions such as loves also having a flatter shape. We find a very similar behavior for engagement with other domains (Supplementary Fig. S4).
@@
In other words: given the PPA $\mathtt{p}$ of a user, and that this user has clicked on URL $u$, what is the probability that the ideology of URL $u$ is in the ideology interval $\mathtt{ib}$? Similarly, we compute the probability of clicking on a URL of domain quality
@@
where $\mathtt{qb}$ denotes a domain quality bin.
Figure 3B shows $P(i(u)\in\mathtt{ib}|\mathtt{p})$ as a heatmap. We see that, while users in the liberal and center classes seem to follow a single-mode distribution, with its center tilted towards negative values for liberals and centered for the center category, conservative users follow a bimodal distribution. We can see one mode at the center and another mode on the far conservative part (upper part in the heatmap). Users with undefined ideology have a very similar profile to center users.

### v3

@@ Abstract
The Facebook Privacy-Protected Full URLs Dataset was released to enable independent, academic research on the impact of Facebook’s platform on society while ensuring user privacy. The dataset has been used in several studies to analyze the relationship between social media engagement and societal issues such as misinformation, polarization, and the quality of consumed news. In this paper, we conduct a comprehensive analysis of the engagement with popular news domains, covering four years from January 2017 to December 2020, with a focus on user engagement metrics related to news URLs in the U.S. By incorporating the ideological alignment and composite score of quality and reliability of news sources, along with users’ political preferences, we construct weighted averages of ideology and quality of news consumption for liberal, conservative, and moderate audiences. This allows us to track the evolution of (i) the ideological gap in news consumption between liberals and conservatives and (ii) the average quality of each group’s news consumption. We identify two major shifts in trends, each tied to engagement changes. In both, the ideological gap widens and news quality declines. However, engagement rises in the first shift but falls in the second. Finally, we contextualize these trends by linking them to two major Facebook News Feed updates. Our findings provide empirical evidence to better understand user behavior and engagement with news and their leaning and reliability during the period covered by the dataset.
@@ 1 Introduction
The interplay between user ideology, news source quality, and online engagement has been a key focus of research in recent years. Studies consistently show that individuals tend to consume ideologically aligned content, with less moderate views often linked to lower-quality news sources Bakshy et al. (2015); Robertson et al. (2023), and more extreme narratives being associated with higher engagement Galeazzi et al. (2024). Extreme conservative audiences, in particular, have been found to exhibit higher levels of ideological segregation González-Bailón et al. (2023); Heseltine (2023). These patterns raise concerns about the relationship between online platforms’ design and the segregation of information consumption, as platform-specific mechanisms can influence the extent of ideological separation. For instance, ideological segregation appears stronger on Facebook than on Reddit, where interactions remain more heterogeneous even among extreme-leaning users Cinelli et al. (2021). Moreover, selective exposure to partisan news sources tends to increase with user activity Cinelli et al. (2020), contributing to the formation of highly clustered news consumption patterns Schmidt et al. (2017).
@@
Building on the literature on algorithmic amplification, recommender systems have been shown to prioritize engagement, often at the expense of exposing users to more polarizing or toxic content, thereby amplifying ideological divides Chavalarias et al. (2024). Regarding the assessment of partisan audience bias in social media, Robertson et al. (2018, 2023) have conducted audits and investigations on user exposure and interactions with content on Google Search, while Bakshy et al. (2015) used US Facebook data on interactions with shared news and friend networks. Contrarily to González-Bailón et al. (2023), both concluded that individual choices played a stronger role in increasing like-minded content, as well as limiting cross-cutting exposure than algorithmic ranking. Additionally, Gentzkow and Shapiro Gentzkow and Shapiro (2011) and Flaxman et al. Flaxman et al. (2016) showed that while ideological filtering existed online, it was partly offset by incidental exposure to diverse viewpoints, particularly through social media These contrasting findings suggest the need for a more comprehensive analysis of user behaviors and algorithmic influences Budak et al. (2024).
@@
The Facebook Privacy Protected Full URLs Dataset Messing et al. (2020) is a large-scale collection of public posts containing URLs and user interaction data on the Facebook platform starting January 2017. The dataset includes counts of engagement at the URL-action level, stratified by limited user attributes such as political page affinity (PPA), age group, and gender. It does not contain, however, information on the volume of unique users in each of these categories or in the dataset in general. While the dataset is unique in its size and granularity, it also includes privacy-protection mechanisms such as injected noise and a minimum-share inclusion threshold that introduce certain limitations. We discuss these in more detail in Section 6.2.
@@
Despite these limitations, researchers have used the Facebook Privacy-Protected Full URLs Dataset for various applications. Our work, akin to Guess et al. (2021) and Bailey et al. (2021), provides a descriptive static analysis of exposure and sharing patterns of news URLs in the US, confirming findings like the conservative skew in low-quality news consumption. Building on this foundation, our work introduces a temporal dimension to investigate ideological segregation and information quality in news consumption. Similar to Buntain et al. (2023), who developed a metric for robust domain ideology scores based on exposure and engagement, we focus on user behavior and content interactions but expand the scope by examining longitudinal trends and dynamic patterns over time.
@@
Prior work on misinformation has shown that platform interventions, such as labeling or demoting low-quality pages, have limited long-term effects Thero and Vincent (2022); Vincent et al. (2022). Bandy and Diakopoulos (2023) show that Facebook amplified low-quality publishers during the 2020 US election, with algorithmic changes impacting publishers broadly, regardless of content quality. We extend this line of inquiry by contextualizing engagement trends within potential algorithmic adjustments Mosseri (2018); Hagey and Horwitz (2021); Narayanan (2023) to examine their implications for ideological segregation on the platform.
@@
Our study contributes specifically to the literature on news audience polarization Mangold and Scharkow (2022), focusing on ideological segregation in engagement on Facebook (i.e. how liberal and conservative users differ in their interactions with news sources over time). Using a four-year longitudinal dataset of Facebook interactions, we characterize engagement trends (e.g., clicks, likes, shares, and comments) across news-related domains. We characterize these patterns in relation to user ideology, domain ideology, and domain quality. By developing a novel metric to assess partisan bias in user consumption patterns, we track the evolution of the ideological gap between news consumed by conservatives and liberals as an indicator for polarization in news diets.
While broader debates on political polarization encompass ideological, affective, and partisan dimensions Baldassarri and Gelman (2008), and findings on the role of online platforms remain mixed Flaxman et al. (2016); Gentzkow and Shapiro (2011); Guess (2021); Boxell et al. (2017); Allcott et al. (2024), our analysis does not aim to assess polarization in society at large. Instead, we examine group-level engagement patterns specific to Facebook, shaped by both user preferences and platform design using a longitudinal dataset.
Additionally, we study changes in the quality of news consumption and how it relates to both user and outlet ideology. To do so, we examine content trustworthiness and integrity using a domain-level quality score based on expert evaluations from Lin et al. Lin et al. (2023), which captures both journalistic standards and a domain’s history of sharing misleading content. While a low-quality score does not equate to the presence of misinformation in every piece of content from that outlet, we interpret low domain quality as indicative of a higher likelihood of such behavior, while avoiding content-level claims. Therefore, our use of the term “misinformation” throughout this work refers to a broad behavioral pattern of spreading false or misleading content, and not the narrower, intent-based definition that distinguishes disinformation.
@@
Our findings provide a detailed view of the dynamic relationship between user behavior, news quality, and ideological segregation in news diets, while situating these changes within the broader context of platform-level adjustments and potential algorithmic modifications.
@@ 2 Data Collection, Processing, and Descriptive Analysis
For the US, these counts are also stratified by the estimated ideological leaning of the users, called Political Page Affinity or PPA, which classifies the audience into five different buckets, scaled from 2-2 (most liberal) to +2+2 (most conservative), with an additional bin (not defined). This metric, based on the model described in Barberá (2015), is computed internally at Facebook. For that reason, we focus on URLs shared primarily in the US and extract engagement counts from users located in the US only.
@@ 2.1 Domain and Data Selection Criteria
For this study, we separate between: (i) news-related domains with reliable annotations for both ideological leaning and quality, and (ii) other domains, that are potentially non-news and include all remaining sites in the Facebook URL dataset, such as entertainment, e-commerce, or low-engagement domains, for which ideological or reliability classifications are not systematically available.
Our selection of news-related domains is conceptually grounded in our goal to examine ideological segregation and information quality in news consumption. As such, we include only domains that (1) are among the top 1% most-viewed domains in the U.S. Facebook dataset (2017–2021), and (2) have both domain-level ideology and quality scores. This filtering is necessary to ensure that each domain in our analysis can be meaningfully positioned along both dimensions.
We construct the set of news domains using the following three sources:
@@
    Top 1% most-viewed domains. We begin with the list of the $2,629$ most-viewed domains Buntain et al. (2023), which account for over 77% of all URL views and over 80% of all URL clicks in the U.S. on Facebook from 2017 to 2020. Given the heavy-tailed distribution of engagement on social media, this threshold is not arbitrary but reflects a pragmatic balance between coverage and signal quality. The domains in this selection range from very high view counts per month (on the order of $10^{11}$) to lower values (of the order of $10^{3}$), reflecting the wide engagement spectrum within the top 1%. Including domains with even lower engagement would introduce substantial noise due to differential privacy constraints, as such domains tend to hover near the dataset’s inclusion threshold and offer a limited usable signal.
@@
    Domain quality scores. We use the domain-level quality scores from Lin et al. (2023), who define a composite quality score based on evaluations from six independent expert rating sources (including NewsGuard [^1] https://www.newsguardtech.com/). These scores reflect a broad range of criteria—such as factual reporting, transparency, editorial standards, and importantly, the frequency with which an outlet has been known to share false or misleading content. The composite score, derived via principal component analysis, captures a latent “domain reputation” dimension that blends aspects of both journalistic professionalism and misinformation-related behaviors. The substantive meaning of the PCA score is not directly interpretable in terms of any one dimension of quality, since it reflects the dominant axis of variation in the rating dataset, which may prioritize some attributes (e.g., factual accuracy) over others (e.g., transparency of sourcing) depending on their variance. We acknowledge that in media and communication studies, the term “quality” often has a restricted normative meaning tied to journalistic professionalism. Our use of the term adopts a broader empirical framing that includes reputational and factual indicators, including a domain’s historical association with false or misleading content. Therefore, a low score may reflect various factors, such as misinformation prevalence or low transparency, without distinguishing between them. Quality scores are available for 1,586 ($\approx 63\%$) of the $2,629$ domains.
  3. 3.
    Ideology scores. We use the domain-level ideology scores from Robertson et al. (2018), who developed a domain-level audience bias metric based on the sharing behavior of registered US voters on Twitter. These scores reflect aggregate partisan preferences in domain sharing and offer a scalable, behavior-based measure of ideological orientation. Compared to manual annotation or content-based scoring, this method avoids rater subjectivity and better captures the political leanings of domains as experienced by users. The metric has been validated against several external sources, showing high agreement with prior audience-based and rater-based measures, and is widely used in polarization and media bias research. These scores are continuous, behavior-based, and include $1,664$ of the $2,629$ domains ($\approx 63\%$).
After intersecting these three sources, we identify 1,231 domains as news-related. This set includes a wide range of outlets: mainstream, local, partisan, centrist, and fringe, and represents a broad and ideologically diverse sample. Among them, approximately 58% have a quality score greater than 0.6, a threshold aligned with NewsGuard’s definition of high trustworthiness Lin et al. (2023). It is important to clarify that the remaining 53% of domains from the original top 1% list are not necessarily excluded news domains. Many are non-news sites (e.g., memes, e-commerce, clickbait farms), and are excluded because they lack ideological or quality annotations. See Table 1 for the number of domains and related URLs for each of the two categories.
@@
As Figure 1 shows, there is a weak negative correlation between domain ideology and domain quality scores (r=0.35,p<0.001r=-0.35,p<0.001), with two clusters of high and medium-high reliability domains located near the ideological center and a concentration of low-quality domains toward the conservative end of the spectrum. Additionally, when considering ideological extremity (i.e., the absolute value of ideology scores), we observe a somewhat stronger negative correlation (r=0.48,p<0.001r=-0.48,p<0.001), indicating that lower quality could be associated with more ideologically extreme domains, regardless of direction.
@@
Figure 1: Selected news domains for this study distributed by Ideology and Quality scores. (bin width ideology =0.094=0.094, quality =0.05=0.05, ticks mark the center of the bin) Positive values of ideology denote a leaning toward conservative ideology, while negative values correspond to liberal leanings. Quality scores are higher when domains have higher quality. The two dimensions correlate negatively (r=0.35,p<0.001r=-0.35,p<0.001).
@@ 2.2 URL Retrieval and Preprocessing
We retrieve URLs through successive monthly queries to the platform, covering the period from January 2017 to December 2020. The dataset is protected by differential privacy Messing et al. (2020); Evans and King (2023), which prevents tracing actions to specific users or URLs. However, while differential privacy safeguards data, it can introduce biases into inference methods if not properly accounted for Evans and King (2023). Additional challenges of this dataset include selection bias caused by engagement thresholds for URL inclusion Allen et al. (2021) and past data integrity issues Allen et al. (2022). An example of a precise query with parameters is detailed in Supplementary Fig. S1, and a discussion of these limitations and our strategies for addressing them can be found in Section 6.2.
@@
To justify our choice of a cutoff period of three months, we calculate the typical lifespan of a URL. Figure 2B shows how engagement decays (on average) over time for the year 2019 (all years behave similarly) for three different types of engagement. We compute the total engagement for a URL within that year and normalize the proportion of engagement for each month, starting from the month in which it was first posted and continuing with subsequent months. The curve illustrates the expected lifespan of a URL’s engagement once it appears in the dataset. We observe that most of the engagement ($\approx 70\%$) occurs during the first two months. We repeated all our analyses using different values for the cutoff period (3-month, 6-month, and 12-month cutoffs) and found that the results for Figures 3 through 7 remained identical across these settings. Because the figures are visually indistinguishable, this data is not shown. This indicates that extending the cutoff period essentially increases the size of the monthly URL set (see 2A), but the overall impact on engagement is negligible. Consequently, we present our results considering engagement counts using a 3-month cutoff.
@@ 3 Aggregate Analysis of Engagement
First, we analyze the distribution of engagement across user PPA categories. Since the dataset is designed to include noisy aggregated interaction counts per PPA (without information about the number of unique users per interaction), we experimented with different baselines and normalization methods. Details on the methods used to compute confidence intervals for engagement counts can be found in the Supplementary Information 1A.
@@
Normalizing by views captures the engagement given the exposure to the content and normalizing by the maximum value allows to compare across different types of engagement, despite their varying volumes. Alternative normalization methods and resulting outcomes are discussed in the Supplementary Information 2.
@@
Figure 3: Aggregate Analysis of Engagement (for popular news domains). In all figures, negative values correspond to more liberal-leaning, and positive values correspond to more conservative-leaning. (A) Distribution of different engagement types across user political page affinity (PPA), normalized relative to view counts (top left). For engagement types other than views, values are further normalized relative to their maximum across the five PPA categories, see Eq. (1). Engagement is unevenly distributed, with liberals having significantly more views than conservatives. The U-shaped pattern of engagement (greater engagement by extreme users) becomes more pronounced from left to right and from top to bottom. Users with undefined PPA exhibit much higher passive engagement (views and clicks) but significantly fewer engagement counts per view and are therefore not represented. (B) Conditional probability of clicking on a URL with a specific ideology given the user PPA. (bin width =0.094=0.094, columns sum one). Liberal and center users follow single-mode distributions of engagement, with liberals leaning towards negative (liberal) domain scores and center users favoring neutral domains. Conservative users display a bimodal distribution, engaging both with neutral and far-right domains. Undefined users exhibit a profile similar to center users. (C) Conditional probability of clicking on a URL with a specific quality given the user PPA. (bin width =0.05=0.05, columns sum one). Liberal users predominantly engage with high-quality domains. Center and conservative users progressively interact with lower-quality domains, with extreme conservatives showing the highest engagement with low-quality domains. This trend is supported by a negative correlation between domain quality and user ideology. See main text for details.
@@
Figure 3A shows the results for popular news domains. Engagement is unevenly distributed, with liberals having significantly more views than conservatives. This suggests that liberals tend to engage more frequently in passive consumption of news-related URL-containing posts, which we find to be opposite to other content (see Supplementary Fig. S3). For clicks, the other type of passive engagement, the pattern shifts slightly to a U-shape, with center users having the lowest number of clicks. As for the active engagement counts, there is a clear U-shaped pattern for hahas, likes, comments, and angers (increasingly pronounced in that same order). Surprisingly, the distribution for shares is the least U-shaped of all, with other reactions such as loves also having a flatter shape. We find a very similar behavior for engagement with other domains (Supplementary Fig. S3).
@@
In other words: given the PPA $\mathtt{p}$ of a user, and that this user has clicked on URL $u$, what is the probability that the ideology of URL $u$, denoted as $i(u)$, is in the ideology interval $\mathtt{ib}$? Similarly, we compute the probability of clicking on a URL of domain quality
@@
where $\mathtt{qb}$ denotes a domain quality bin and $q(u)$ the domain quality of URL $u$ (see Section 4.2 for a formal definition of these variables).
Figure 3B shows $P(i(u)\in\mathtt{ib}|\mathtt{p})$ as a heatmap. We see that, while users in the liberal and center classes seem to follow a single-mode distribution, with its center tilted towards negative values for liberals and centered for the center category, conservative users follow a bimodal distribution. We can see one mode at the center and another mode on the far conservative part (upper part in the heatmap).
@@
Users with undefined ideology (referred to as “Not defined” in the figures) show engagement patterns similar to center users. This is supported by the Total Variation distance between the distribution of “Not defined” users and each ideological group: {-2: 0.33, -1: 0.26, 0: 0.06, 1: 0.31, 2: 0.41} for ideology-based engagement (Figure 3B), and {-2: 0.10, -1: 0.09, 0: 0.08, 1: 0.22, 2: 0.31} for quality-based engagement (Figure 3C). In both cases, “Not defined” users are most similar to center users in their engagement patterns.